hhanan:

What is the next term in the sequence 2,-2/3, 2/9,-2/27...

Hero:

Hint: \(3^0 = 1\) \(3^1 = 3\) \(3^2 = 9\) \(3^3 = 27\)

hhanan:

thank you

jhonyy9:

so bc. you need getting from 2 in the first step -2/3 ,what you think what will be a general formula of this given number line ?

Rahima:

2+4=8, 3+8=17, 5-5=20, 6-5=?

Rahima:

Plz solve it

jhonyy9:

2*2=4+4=8 3*3=9+8=17 5*5=25-5=20 6*6=36-5=31 have you get the way step by step ?